Let (X,Y) be a two-dimensional continuous random variable with joint P.d.f







π‘“π‘‹π‘Œ







(π‘₯,𝑦) = {







2 , 0 < π‘₯ < 𝑦 < 1







0 , π‘œπ‘‘β„Žπ‘’π‘Ÿπ‘€π‘–π‘ π‘’







Find the marginal distribution for X and Y ?

1
Expert's answer
2022-03-29T08:20:39-0400

"f_X\left( x \right) =\int{f\left( x,y \right) dy}=\int_x^1{2dy}=2\left( 1-x \right) ,x\in \left[ 0,1 \right] \\f_Y\left( y \right) =\int{f\left( x,y \right) dx}=\int_0^y{2dx}=2y,y\in \left[ 0,1 \right]"